After there are several waves in the simulation pause it and use the mouse to find...

  1. After there are several waves in the simulation pause it and use the mouse to find the wavelength (distance between two successive crests). What are the wavelength and speed of the wave (wavelength/period)?
Answer #1

the distance between two crests = wavelength

so wavelength = 0.125-0.025 = 0.1 m

and time period = 1/3 =0.333 sec

therefore, speed = wavelength/time period = 0.3 m/s